An air-standard cycle is executed in a closed system and is composed of the following four processes:

1-2 Isentropic compression from 100 kPa and 27°C to 1 MPa

2-3 P = constant heat addition in the amount of 2800 kJ/kg

3-4 v = constant heat rejection to 100 kPa

4-1 P = constant heat rejection to initial state

(a) Show the cycle on P-v and T-s diagrams.

(b) Calculate the maximum temperature in the cycle.

(c) Determine the thermal efficiency.

Assume constant specific heats at room temperature.


1
Expert's answer
2021-11-29T19:08:36-0500

Given ,

  • = 100 KPa
  • = 27 C = 300 K.
  • = P = 1 MPa =1000 KPa.
  • 23 = 2800 KJ/kg.
  • = V .
  • = 100 KPa.

a)


b) For process 1 - 2

"\frac{T_2}{T_1}=(\frac{P_2}{P_1})^{\frac{γ-1}{γ}} \\

T_2=300(\frac{1000}{100})^{\frac{1.4-1}{1.4}}"

T2=579.2K


For heat addition 2-3 -

"Q_23=C_p(T_3 -T_2)"

2800=1.005(T3 - 579.2)

T3 =3365.26K=3092.25C


c) For process 3-4 

"\frac{T_4}{T_3}=\frac{P_4}{P_3} \\

T_4=3365.26 \times \frac{100}{1000} \\

T_4=336.526K"

Total heat rejected in the cycle -

"Q_r=Q_34 +Q_41 \\

Q_r=C_v (T_3 -T-4) + C_p (T_4 - T_1) \\

Q_r= 0.718(3365.26 - 579.2) - 1.005(336.526 - 300) \\

Q_r=2037.009 Kj/kg"



Now,

Work done in the cycle will be -

W=Qsupplied − Qrejected

W=2800−2037.099

W=762.901KJ/kg.


Efficiency of the cycle will be -

"n = \frac{Workdone}{heatsuppied} \\

n = \frac{762.901}{2800} \\

n=0.2724= 27.24" %